Find the escape speed from a planet with a mass of 3×10^24 kg and a diameter of 2×10^6 m. Use 2⁄3×10^-10 N·m2/kg2 for the gravitational constant.
Answer #1

Solution)

Escape speed,

V= sqrt( 2GM/R) = sqrt((2*2/3*10^-10*3*10^24)/10^6) =20,000 m/s (Ans)

=======